# Enhanced sensitivity of sub-THz thermomechanical bolometers exploiting vibrational nonlinearity

### v2

@@ III Spectral transduction
Moreover, the PyC device shows about an order of magnitude improvement with respect to the Au device. This originates from the different absorbance of the two materials at 140 GHz; interestingly, comparing the best NEP values in the two devices, opportunely rescaled to account for the different slopes, gives an absorption-induced enhancement of $\sim$3.74, in good agreement with the scaling of the static responsivities estimated as $R_{s}^{PyC}\sim$1042 kHz/W and $R_{s}^{Au}\sim$278 kHz/W, giving $R_{s}^{PyC}/R_{s}^{Au}=3.75$. From this observation, we can estimate the absorbance of the metallic layer, starting from an experimentally measured PyC film absorbance of 40% in the far-infrared range [49]. The obtained Cr/Au absorbance of $\sim$10is compatible with a simple multilayer simulation using experimentally estimated refractive indices [51], which returns a numerical value of $4\pm 0.5\%$ for a continuous film, whereas we expect a further enhancement due to the layer inhomogeneity. The best NEP we obtained reaches 0.2 $\mathrm{pW/\sqrt{Hz}}$, entering the range of state-of-the-art diode [52, 53] and transistor [54, 55, 56] detectors, at the expense of a slower operating speed and reduced dynamic range. As a more quantitative example, with the driving condition of Figure 3, the PyC device negative derivative peak has a linewidth of about 150 Hz, limiting its use to weak signals with power less than roughly 140 nW, as estimated by considering the spectral static device responsivity.
@@
A further increase of the driving voltage leads the system to highly nonlinear regimes which, in the case of Duffing nonlinearities, can eventually reach multistable or chaotic motion [39]. As an example, Figure 4 (a-b) reports the comparison of an asymmetric, single-solution resonance (red curve) and a regime past the bifurcation point (blue curve) for both devices under investigation. The multistable regime can be characterized by abrupt jumps resulting from the switching between two dynamical solutions, which can be clearly identified in both the PyC device and Au device. The onset of multistability essentially depends on the vibrational amplitude, since the Duffing nonlinearity enters the equation of motion with a cubic displacement term [36]. The bifurcation point from a single to multiple dynamical solution has been taken as a common reference to compare both devices.
@@
One can think of the switching as a “line with infinite slope”, which looks appealing for transduction detection. Unfortunately, in this case the derivative tends to a $\delta$-function, as can be seen from the ones reported in Figure 4 (c-d), respectively. Evaluating the figures (and the insets for the full range of the $y$ axis), one can expect an extreme enhancement of device performance in terms of responsivity, which comes along with a vanishing dynamic range.
@@
Since it operates at a single frequency point, this particular regime of operation does not grant a use for the device as an imaging or power meter detector; conversely it looks promising for a threshold-based detector, where the discontinuity in the read-out can be triggered by specific events over a certain limit (i.e. single photons, light pulse detection, temperature change, mass loading). In a scheme similar to the superconducting optical detectors [20] or bifurcation-based sensors [25, 26, 27, 28, 29] in other nonlinear systems, this operating regime adds to the potential and possibilities lying in thermomechanical bolometer devices.
@@ IV Conclusion
While increasing the quality factor in resonant detectors generally enhances device performance, consistently achieving high values can be challenging, especially in the context of mass production for real-world applications. In this work, we presented an alternative approach to locally achieve high responsivity and low NEP in transduction detection experiments by exploiting device nonlinearities.
Through careful characterization of thermomechanical bolometers, we demonstrated that the device NEP under 140 GHz illumination is strongly dependent on the driving strength and transduction frequencies. In the optimal operating range, we achieved a sub-pW/$\sqrt{\mathrm{Hz}}$ NEP for a TMB employing pyrolitic carbon as an absorbing layer, which exhibited an impressive 40% absorbance in just an 18 nm thick layer [49].
This carbon-based material also offers additional advantages: not only it is chemically inert, but its thickness and thus its AC conductivity and absorption capability can be conveniently controlled by tuning the synthesis conditions [57].
With a 40 Hz operating speed, our device offers a powerful tool for power detection of low-intensity signals.
Given  that the increase of the trampoline temperature exposed to electromagnetic radiation is based on broad band absorption in a thin conductive layer (PyC or Au) we expect our approach to provide as good performance for detection of radiation up to 10 THz.
Additionally, by further increasing the driving strength and entering a highly nonlinear regime, we propose leveraging the same platform for threshold signal detection. In this regime, the system can transition between stable solutions in response to external perturbations of sufficient magnitude, similar to superconducting or bifurcation-based detectors, opening up further possibilities for our technology.
Acknowledgements
The authors gratefully acknowledge funding from ATTRACT, a European Union’s Horizon 2020 research and innovation project under grant agreement No. 101004462 (H-cube project), from European Union - Next Generation EU under the Italian National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.3, CUP D43C22003080001, partnership on “Telecommunications of the Future” (PE00000001 - program “RESTART”) and from MUR through the project PRIN 2022 TRUST. This work is supported by the Academy of Finland via Flagship Programme Photonics Research and Innovation (PREIN), decision 320166, Horizon 2020 RISE CHARTIST project 101007896, Horizon 2020 RISE TERASSE project 823878.

### v3

@@ III Spectral transduction
The NEP evaluated by setting $f_{D}$ at the maximum and minimum slopes of both devices is reported in Figure 3 (e) as a function of the modulation frequency of the 140 GHz source. All NEP curves remain relatively constant as a function of $f_{M}$, consistent with the results of Figure 2 (d), and show an overall low value which strongly depends of the chosen transduction frequency. As expected from Eqs. (3) and (2), the NEP is inversely proportional to the derivative, resulting in net reductions of about a factor of 117 and 37.5, respectively, in good agreement with what one would expect from the results of Fig. 3 (b) and (d). Note that the NEP reductions are obtained within the same mechanical device, and obtained just by changing $f_{D}$. Moreover, the PyC device shows about an order of magnitude improvement with respect to the Au device which originates from the different absorbance of the two materials at 140 GHz. Interestingly, comparing the best NEP values in the two devices, appropriately re-scaled to account for the different slopes, gives an absorption-induced enhancement of $\sim$7.94, in good agreement with the ratio of static responsivities, evaluated as $R_{s}^{PyC}\sim$497 MHz/W and $R_{s}^{Au}\sim$72.6 MHz/W, giving $R_{s}^{PyC}/R_{s}^{Au}=6.85$. From this observation, we can estimate the absorbance of the metallic layer, starting from an experimentally measured PyC film absorbance of $\sim$40in the far-infrared range [49]. The obtained Cr/Au absorbance of $\sim$10is compatible with multilayer simulations using the experimentally estimated refractive indices of Si3N4 and metals [51], which returns a numerical value of $4\pm 0.5\%$ for a continuous film, whereas we can assume a further enhancement due to the layer inhomogeneity.
The best NEP we obtained is about 30 $\mathrm{pW/\sqrt{Hz}}$ for the PyC device. As reported in Figure 3(e), we still have room towards the fundamental NEP limit for a detector operating at a temperature $\mathrm{T}$, set by thermal fluctuations in the system and defined as [52]:
$NEP^{*}=\sqrt{16k_{B}T^{5}\sigma A}$ (4)
where $k_{B}$ is the Boltzmann constant, $\sigma$ is the Stefan–Boltzmann constant and $\mathrm{A}$ is the absorber area. Even if some thermal detectors shows exceptional performances and metrics close to the thermal limit in the mid-infrared range, they rely on all-optical probes, which are known to inject less noise to the system with respect to the all-electrical ones employed here [50]. Conversely, all-electrical read-out is better suitable for portability and integration and our NEP compares well with some of the best commercial devices in the sub-THz range, which have NEPs around 10 $\mathrm{pW/\sqrt{Hz}}$ [53]. The main drawback of our technique lies in the reduced bandwidth where we find very large derivatives. This directly translates into a reduction of the detector dynamic range (i.e. the ratio between the maximum and minimum detectable power). As a quantitative example, with the driving condition of Figure 3, the PyC device negative derivative peak has a linewidth of about 50 Hz, limiting its use to weak signals with power less than roughly 100 nW, as estimated by considering the spectral static device responsivity.
@@
A further increase of the driving voltage leads the system to highly nonlinear regimes which, in the case of Duffing nonlinearities, can eventually reach multistable or chaotic motion [39]. As an example, Figure 4 (a-b) reports the comparison of an asymmetric, single-solution resonance (red curve) and a regime past the bifurcation point (blue curve) for both devices under investigation. The multistable regime is characterized by abrupt jumps resulting from switching between two stable solutions; the onset of multistability essentially depends on the vibrational amplitude, since the Duffing nonlinearity enters the equation of motion with a cubic displacement term [36]. The bifurcation point from a single to multiple dynamical solutions has been taken as a common reference to compare both devices.
@@
One can think of the switching as a “line with infinite slope”, which looks appealing for transduction detection. Unfortunately, in this case the derivative tends to a $\delta$-function, as can be seen from the normalized numerical derivates reported in Figure 4 (c-d), respectively. Evaluating the figures (and the insets for the full range of the $y$ axis), one can expect an extreme enhancement of device performance in terms of responsivity, which comes along with a vanishing dynamic range.
@@
Since diverging derivatives can be found at a single frequency point, this particular regime of operation does not allow utilizing the device as an intensity detector; conversely it looks promising for a threshold-based sensor, where the discontinuity in the read-out can be triggered by specific events over a certain limit (i.e. single photons, light pulse detection, temperature change, mass loading). In a scheme similar to the superconducting optical detectors [20] or bifurcation-based sensors [25, 26, 27, 28, 29] in other nonlinear systems, this operating regime adds to the potential and possibilities lying in thermomechanical bolometer devices.
@@ IV Conclusion
While increasing the quality factor in resonant detectors generally enhances device performance, consistently achieving high values can be challenging, especially in the context of mass production for real-world applications. In this work, we presented an alternative approach to locally achieve high responsivity and low NEP in transduction detection experiments by exploiting device nonlinearities. Through careful characterization of thermomechanical bolometers, we demonstrated that the device NEP under 140 GHz illumination is strongly dependent on the driving strength and transduction frequencies. In the optimal operating range, we achieved a NEP of 30 $\mathrm{pW/\sqrt{Hz}}$ for a TMB employing pyrolitic carbon as an absorbing layer, which exhibits a 40% absorbance in the sub-THz range under investigation [49]. Additionally, by further increasing the driving strength and entering a highly nonlinear regime, we propose leveraging the same platform for threshold signal detection. In this regime, the system can transition between stable solutions in response to external perturbations of sufficient magnitude, with a behaviour similar to superconducting or bifurcation-based detectors, opening up further possibilities for our technology.
V Acknowledgment
The authors gratefully acknowledge funding from ATTRACT, a European Union’s Horizon 2020 research and innovation project under grant agreement No. 101004462 (H-cube project), from European Union - Next Generation EU under the Italian National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.3, CUP D43C22003080001, partnership on “Telecommunications of the Future” (PE00000001 - program “RESTART”) and from MUR through the project PRIN 2022 TRUST. This work is supported by the Academy of Finland via Flagship Programme Photonics Research and Innovation (PREIN), decision 368653, Horizon 2020 RISE CHARTIST project 101007896, Horizon 2020 RISE TERASSE project 823878.